OpenStudy (anonymous):

What is 3m + 5 = 8m - 105

OpenStudy (anonymous):

8m - 3m = 5 + 105 Can you solve it now???

OpenStudy (anonymous):

I'm suppose to do 3m + 5 = 8m - 105, I got 11m + 5 = -105?

OpenStudy (anonymous):

after that I got...OH lol nvm tyvm you two :)

OpenStudy (anonymous):

Whoops, made a careless mistake O.o

OpenStudy (anonymous):

See, when you bring 8m from right hand side to left hand side, it becomes negative on coming to left hand side.. 3m - 8m +5 = -105 Getting my point or not??

OpenStudy (anonymous):

Yup, I got it...Math is just the best.

OpenStudy (anonymous):

So, what you got??

OpenStudy (anonymous):

110/11= 10

OpenStudy (anonymous):

Sorry friend you are wrong....

OpenStudy (anonymous):

On Subtracting 3 from 8, what will you get??